assume that the random variable X is normally distributed, with mean m = 50 and standard deviation s = 7. Find each indicated percentile for X.
The 90th percentile?

Answer by ewatrrr(24785)

90% then z = invNorm(.90) = 1.2816
z = 1.2816 = (X-50)/7
7(1.2816) + 50 = 8.97 + 50 = X = 59